A fluid undergoes a reversible adiabatic compression from 0.5 MPa, 0.2m^3 to 0.05m^3 according to the law PV^1.3=constant. Determine the change in enthalpy, internal energy and heat transfer and work transfer during the process.
1
Expert's answer
2019-04-18T10:00:11-0400

Adiabatic process is a process in which a thermodynamic system does not exchange heat or substance with its surroundings. Hence, in the given adiabatic compression, the heat transfer "Q" is equal to zero.


The first quantity we can determine is the work done by the system on its surroundings. It is calculated by taking the following integral:

"W = \int_{V_1}^{V_2} p dV \, ,"

where "p" is pressure, and "V_1 = 0.2\, \text{m}^3" and "V_2 = 0.05\, \text{m}^3" is the initial and final volume, respectively. The pressure as a function of volume is provided by the conditions of the problem: "p V^{1.3} = \text{const} = p_1 V_1^{1.3}", where "p_1 = 0.5\, \text{MPa} = 0.5 \times 10^6\, \text{N}/\text{m}^2" is the initial pressure. Hence,

"p = p_1 \left( \frac{V_1}{V} \right)^{1.3} \, ."

Substituting this into the above integral and performing the elementary integration, we obtain

"W = \frac{p_1 V_1}{0.3} \left[ 1 - \left( \frac{V_1}{V_2} \right)^{0.3} \right] \approx - 1.72 \times 10^5\, \text{J}\, ,"

where we have taken into account that "\text{N} \cdot \text{m} = \text{J}". Thus, the positive work is actually done by the surroundings on the system.


The change "\Delta U" in the internal energy of the system is determined from the first law of thermodynamics:

"\Delta U = Q - W = - W = 1.72 \times 10^5\, \text{J} \, ,"

since "Q = 0".


The enthalpy of a system is defined as "H = U + p V". The change in enthalpy in the given process is, therefore,

"\Delta H = \Delta U + p_2 V_2 - p_1 V_1 \, ."


To find the product "p_2 V_2", we use the equation "p_2 V_2^{1.3} = p_1 V_1^{1.3}", whence "p_2 V_2 = p_1 V_1 \left( V_1 / V_2 \right)^{0.3}". Eventually,

"\Delta H = \Delta U + p_1 V_1 \left[ \left( \frac{V_1}{V_2} \right)^{0.3} - 1 \right] \\ {} = - W - 0.3\, W = - 1.3 \, W \approx 2.24 \times 10^5\, J \, ."